40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1426     Entry time: Wed Mar 25 04:18:28 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Tuesday Locking 
After the new PO_DC PD was installed, I tweaked several gains to make the locking scripts work right.
First of all, I increased the gain of PD12 (PD12_I is SPOB) by a factor of 1.4 to compensate for the power decrease
by the insertion of the BS. SPOB is used by the PRM alignment script. I was too lazy to modify the scripts.

Then I optimized the SRC DD signal which is taken from the POB.

I also had to do some gain adjustments for the CARM loop.

The attachment (AO path open loop TF) shows a depressing fact that the 3.8kHz peak is still there with the new PO_DC PD. So it was not a problem of the SPOB PD.
Next, I will check the cross over frequencies of the PZT and PC paths in the FSS and the VCO/MCL cross over.
Attachment 1: AO-Loop-p9.png  60 kB  | Hide | Hide all
AO-Loop-p9.png
ELOG V3.1.3-